Shift report 20030210 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Feb 10 2003 - 23:54:12 EST

  • Next message: Apache: "Shift report 20030211 00:00-08:00"
    ------------------------------------------------------------
    Date: Mon Feb 10 23:54:11 2003, Shift: 16:00-24:00
    Supervisor: ZM, with: MG
    Log book#14 pages 9 - 9
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine status ok during all shift. Very low <br>
    background. MCR used some colliders for the 2 beams. <br>
    
    2. Very good and long run store. <br>
        
    -Ended run 7861 from previous shift. 50 k stored <br>
    for tr 6, 250 stored for tr 3. <br>
    -Now running 7861. Until now 158 k stored for tr 3, 61 k <br>
    stored for tr 6.<br>
    
    3. No detector problems. <br>
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Feb 10 2003 - 23:58:08 EST